OpenStudy (anonymous):

What is the solution to the following system of equations? y = x2 + x + 8 y = x2 + 2x + 5 (2, 1) (5, 8) (3, 20) (2, 40)

OpenStudy (anonymous):

\[at X ^{2}+X+8=X ^{2}+2X+5\] then "X"=8-5=3 ;by subistitution "Y"=20